# Linguistic Neuron Overlap Patterns to Facilitate Cross-lingual Transfer on Low-resource Languages

### v1

@@ Abstract
The current Large Language Models (LLMs) face significant challenges in improving performance on low-resource languages and urgently need data-efficient methods without costly fine-tuning. From the perspective of language-bridge, we propose BridgeX-ICL, a simple yet effective method to improve zero-shot Cross-lingual In-Context Learning (X-ICL) for low-resource languages. Unlike existing works focusing on language-specific neurons, BridgeX-ICL explores whether sharing neurons can improve cross-lingual performance in LLMs or not. We construct neuron probe data from the ground-truth MUSE bilingual dictionaries, and define a subset of language overlap neurons accordingly, to ensure full activation of these anchored neurons. Subsequently, we propose an HSIC-based metric to quantify LLMs-internal linguistic spectrum based on overlap neurons, which guides optimal bridge selection. The experiments conducted on $2$ cross-lingual tasks and $15$ language pairs from $7$ diverse families (covering both high-low and moderate-low pairs) validate the effectiveness of BridgeX-ICL and offer empirical insights into the underlying multilingual mechanisms of LLMs.
@@ 1 Introduction
Although Large Language Models (LLMs) have demonstrated impressive multilingual capacities, there is still significant space for improving the performance on low-resource languages Huang et al. (2024); Nazi et al. (2025). To address this issue, especially avoiding costly post-training Muller et al. (2021); Yong et al. (2023), it is critical to fully investigate the multilingual understanding and transferring ability in LLMs.
@@
Recent research has increasingly focused on data-efficient methods, particularly Cross-lingual In-Context Learning (X-ICL) Winata et al. (2021); Tanwar et al. (2023); Nazi et al. (2025); Cahyawijaya et al. (2024), which surprisingly works well on low-resource languages, primarily due to LLMs are in-context low-resource language learners Brown et al. (2020b); Cahyawijaya et al. (2024). For instance, in the Arabic-to-Hebrew Bilingual Lexicon Induction (BLI) task, the zero-shot baseline accuracy in LLaMA 3 is 47.0%. However, simply specifying English as a bridge language in a zero-shot setting boosts accuracy to 64.5%, which significantly outperforms even the two-shot X-ICL. This observation motivates us to further explore: How can we improve cross-lingual capabilities of LLMs on low-resource languages by selecting an optimal bridge language in X-ICL? Should the selection be purely data-driven, favoring high-resource bridge languages Vulic et al. (2020)? Or can human linguistic knowledge, such as language genealogy, or established evolutionary taxonomies, offer a more effective alternative Stanczak et al. (2022); Wang et al. (2024)?
@@
To fully investigate this issue from a systematic perspective, we leverage linguistic neuron Tang et al. (2024) to guide optimal bridge language selection in X-ICL. However, there are two limitations when applying neuron-based interpretation Cao et al. (2024); Tang et al. (2024); Liu et al. (2024) on low-resource languages:
@@
$\bullet$ Inaccurate Neuron Activation. Current work often relies on multilingual corpora like Wikipedia Foundation (2024) to probe internal-neurons, without verifying whether LLMs truly understand the input. This may lead to unreliable neuron activations, particularly for low-resource languages. When LLMs poorly understand the probe input, they may instead activate neurons largely for processing unfamiliar or noisy input.
@@
$\bullet$ Lacking guidance for cross-lingual transfer. Current work focuses on analyzing the distribution of neurons Stanczak et al. (2022); Tang et al. (2024); Liu et al. (2024), yet how to leverage internal neurons to enhance cross-lingual transfer remains underexplored. Recent work argues that language-specific neurons do not facilitate cross-lingual transfer Mondal et al. (2025). This raises a critical question: Whether sharing neurons can improve cross-lingual transfer in LLMs? This exploration is also important to transfer language neuron research to actionable strategies for enhancing LLMs’ multilinguality.
@@
Motivated by this, we propose a simple yet effective bridge method, BridgeX-ICL, to improve LLMs’ cross-lingual capabilities, especially on low-resource languages. To address the inaccurate activation issue, we construct probe data by leveraging the ground-truth bilingual lexicon MUSE Conneau et al. (2017). We then collect bilingual word pairs that LLMs can translate accurately and prompt them to LLMs using answer generation in bi-directions. To address the cross-lingual guidance issue, we first explore overlap neurons’ features and their impact on cross-lingual transfer, and then propose a bridge selection strategy based on HSIC Gretton et al. (2005). Furthermore, we measure the linguistic spectrum in LLMs based on overlap neurons and compare it with human language genealogy from Glottolog Trees Hammarström et al. (2023). We conduct extensive experiments on $2$ cross-lingual tasks and $15$ language pairs from $7$ diverse families. Our main contributions and findings are as follows:
@@
  • We construct neuron probe data and use them to fully activate the anchored overlap neurons. We also propose a HSIC-based metric to quantify the similarity between overlapping-neurons and specific neurons for making optimal bridge selection in X-ICL.
  • We validate the efficacy and generalization of BridgeX-ICL on 2 cross-lingual tasks and $15$ language pairs. Here are empirical findings: 1) Strong neural overlaps align with human linguistic taxonomy within language families, but do not consistently hold across families. 2) Overlap neurons embody shared semantic information no matter among languages within or across families. 3) BridgeX-ICL improves the performance on 2 cross-lingual tasks across 15 language pairs by an average of 6.02% and 5.25% over zero-shot baselines.
@@ 2.1 Cross-lingual In-context Learning
LLMs face significant challenges when applied to low-resource languages Costa-jussà et al. (2022); Muennighoff et al. (2023); Huang et al. (2024), mainly due to insufficient training data and the curse of multilinguality Conneau et al. (2020). To address these issues without updating model parameters, Cross-lingual In-context Learning (X-ICL), an extension of in-context learning (ICL), has recently gained attention Brown et al. (2020a). Prior studies Winata et al. (2021); Tanwar et al. (2023); Nazi et al. (2025); Cahyawijaya et al. (2024) have demonstrated that LLMs act as effective few-shot multilingual learners, with few-shot ICL even outperforming fine-tuned language-specific models on several tasks Winata et al. (2021). However, few-shot X-ICL’s performance is highly dependent on the context and the selection of examples, especially for unconventional or ambiguous languages Philippy et al. (2023); Nazi et al. (2025). Consequently, existing research mainly focuses on optimizing few-shot example selection. To the best of our knowledge, we are the first to explore X-ICL explicitly from the perspective of leveraging language bridges.
@@ 2.2 Linguistic Neuron in LLMs
Recent research Stanczak et al. (2022); Tang et al. (2024); Liu et al. (2024); Cao et al. (2024); Wang et al. (2024) has revealed that language-related neurons exist in FFN layers of transformer architecture. Deactivating these neurons will play a vital impact on LLMs’ multilingual capacities. Beyond uncovering multilingual mechanisms, some research has gone to explore the neuron pattern across languages Wang et al. (2024); Stanczak et al. (2022) and its impact on cross-lingual performance Mondal et al. (2025); Zhang et al. (2025). Specifically, Wang et al. (2024) observed that similar languages may not exhibit significant neuron sharing in LLMs like BLOOM, suggesting that neuron sharing does not fully align with language similarity. Furthermore, recent work argues that language-specific neurons do not facilitate cross-lingual transfer Mondal et al. (2025). This raises a critical question: Whether sharing neurons can improve cross-lingual performance in LLMs. Motivated by these findings, we aim to further investigate LLM-internal neuron sharing across languages and its impact. In particular, we define a subset of language-overlapping neurons and explore whether they can serve as internal bridges to support cross-lingual inference.
@@ 3.1 Task Statement
Given language set $\mathcal{L}=\{L_{1},...L_{|\mathcal{L}|}\}$, this work aims to measure the linguistic genealogy learned by LLMs from language-overlapping neurons and then use the quantified linguistic similarity to guide the bridge language selection in X-ICL.
Figure 1 depicts three main steps of our approach: ① Neuron probe data construction; ② Language neurons and their overlappings detection; ③ Bridge language selection based on a modified HSIC dependency estimation to measure linguistic distance, selecting $L_{*}$ from the candidate set $L_{\text{candidate}}$ to facilitate X-ICL from source language $L_{s}$ to target language $L_{t}$.
@@ 3.2 Probe Data Construction
We utilize two types of probe data for language neuron identification and optimal bridge selection. The former focuses on language-specific neurons and can use existing multilingual corpora. In our work, we adopt FLORES+ NLLB Team et al. (2024), a high-quality parallel corpus released by Meta, and combine its devtest and test sets to obtain 2,000 parallel sentences for each language.
Bridge selection for X-ICL needs to consider both language-specific neurons and those involved in cross-lingual tasks. Inspired by findings in task-specific neurons Song et al. (2024), we hypothesize certain neurons are associated with cross-lingual transfer whose manipulation and measurement should not rely on feeding monolingual input. Therefore, we construct probe data by leveraging bilingual word translations.
Prompt Design We collect $d$ (i.e., $100$) word pairs that LLMs can translate accurately per language pair. We prompt these word pairs to LLMs in both directions of $L_{1}$-$L_{2}$ and $L_{2}$-$L_{1}$, ensuring neurons linked to $L_{1}$ and $L_{2}$ are fully activated. Instead of feeding word pairs directly, we prompt LLMs to generate translations, which guarantees accurate and pronounced neuron activation. Examples of probe data for $3$ language pairs are shown below.
@@ 3.3.1 Neurons in LLMs
Neurons identification is based on Tang et al. (2024), which hypothesizes that language neurons are mainly located in the Feed-Forward Network (FFN) layers. Given the transformation in the $i$-th layer:
@@
where $\tilde{\boldsymbol{h}_{i}}$ is the hidden state input to the $i$-th layer and $\sigma(\cdot)$ denotes the activation function. $\boldsymbol{W}_{1}^{i}\in\mathbb{R}^{d\times N}$ and $\boldsymbol{W}_{2}^{i}\in\mathbb{R}^{N\times d}$ are the learned parameters. Here, a neuron is defined as a linear transformation of a single column in $\boldsymbol{W}_{1}^{i}$ and there are $N$ neurons in each layer. The activation value of the $j$-th neuron is $\sigma(\tilde{\boldsymbol{h}_{i}}\boldsymbol{W}_{1}^{i})_{j}$. If this value exceeds 0, the neuron is considered activated.
@@ 3.3.2 Overlap Neuron Identification
First, we identify neurons $\mathcal{T}_{k}$ associated with specific languages $L_{k}$. Unlike existing work Wang et al. (2024); Mondal et al. (2025) using LAPE Tang et al. (2024) to identify neurons with high activation probability for one language but low for others, which is less effective to detect neuron relationships across languages, we identify neurons $\mathcal{T}_{k}$ for each language $L_{k}$ based their activation frequency. Let $f_{k,j}$ denote the activation frequency of neuron $n_{j}$ when processing tokens of language $L_{k}$. Neurons with the top $\tau\cdot N$ activation frequencies are selected into $\mathcal{T}_{k}$ based on a threshold $\tau$.
Overlap Neuron Definition. For languages $L_{u}$, $L_{v}$, given their associated neurons $\mathcal{T}_{u}$ and $\mathcal{T}_{v}$, the overlap neurons are the interaction of $\mathcal{T}_{u}$ and $\mathcal{T}_{v}$. For the $i$-th layer, we have $\mathcal{T}_{u,v}(i)=\mathcal{T}_{u}(i)\cap\mathcal{T}_{v}(i)$.
Linguistic Neuron Similarity. We measure the linguistic similarity between $L_{u}$ and $L_{v}$ based on their overlap neurons’ activation frequency as:
@@
where $f_{u,j}\in\boldsymbol{f_{u}}$ and $f_{v,j}\in\boldsymbol{f_{v}}$ denote the activation frequency of the $j$-th neuron in $\mathcal{T}_{u,v}$ when processing tokens from $L_{u}$ and $L_{v}$, respectively.
3.3.3 Overlap Neurons Patterns
Second, we use the constructed probe data to explore overlap neurons’ features and their generalized impact on cross-lingual transfer so that we can utilize them to guide bridge language selection. We make two observations:
@@ 3.3.3 Overlap Neurons Patterns
Figure 3: Layer-wise latent embeddings projected with MDS in French-Swahili translation. A rainbow-colored path traces the latent embeddings across 32 layers. The predicted Swahili tokens are in green and their correct English tokens in orange.
$\bullet$ Similar languages share more neurons than distant ones. For example, Arabic-Hebrew within the same language family has more overlapping-neurons than Arabic-Swahili across families, presented in Figure 2, This suggests the potential of neural overlap to measure language distance.
$\bullet$ Overlap neurons concentrate in middle-layers and final-layers and serve distinct roles of semantic understanding and language coding for next-token prediction. This is also evidenced by neurons deactivation in Figure 7. Neurons in final layers are task-related and are responsible for cross-lingual generation. To examine whether middle-neurons handle semantic understanding, we employ a technique called logit lens Nostalgebraist (2020) to visualize the latent embeddings passing through neuron layers. We test in French-Swahili translation and select 60 word pairs that LLaMA 3 can translate accurately. After recording the model’s latent embeddings at each layer for next-token prediction, we use classical multidimensional scaling (MDS) to embed them in a 2D space, presented in Figure 3. The embedding trajectory is marked in a rainbow-colored path (e.g., red = layers 1-4, violet = layers 25-32). We can observe French inputs and the correct English next tokens cluster in middle layers, suggesting LLMs rely on the knowledge in high-resource languages like English to perform cross-lingual reasoning. Neurons in middle-layers should be prioritized over final-layers when measuring language similarity.

### v2

@@ Abstract
The current Large Language Models (LLMs) face significant challenges in improving their performance on low-resource languages and urgently need data-efficient methods without costly fine-tuning. From the perspective of language-bridge, we propose a simple yet effective method, namely BridgeX-ICL, to improve the zero-shot Cross-lingual In-Context Learning (X-ICL) for low-resource languages. Unlike existing works focusing on language-specific neurons, BridgeX-ICL explores whether sharing neurons can improve cross-lingual performance in LLMs. We construct neuron probe data from the ground-truth MUSE bilingual dictionaries, and define a subset of language overlap neurons accordingly to ensure full activation of these anchored neurons. Subsequently, we propose an HSIC-based metric to quantify LLMs’ internal linguistic spectrum based on overlapping neurons, guiding optimal bridge selection. The experiments conducted on $4$ cross-lingual tasks and $15$ language pairs from $7$ diverse families, covering both high-low and moderate-low pairs, validate the effectiveness of BridgeX-ICL and offer empirical insights into the underlying multilingual mechanisms of LLMs. The code is publicly available at https://github.com/xuyuemei/BridgeX-ICL.
Linguistic Neuron Overlap Patterns to Facilitate Cross-lingual Transfer on Low-resource Languages
Yuemei Xu1, Kexin Xu1, Jian Zhou1, Ling Hu1, Lin Gui2 1 School of Information Science and Technology, Beijing Foreign Studies University 2 Department of Informatics, King’s College London {xuyuemei, xukexin, bwzj, huling}@bfsu.edu.cn lin.1.gui@kcl.ac.uk
@@ 1 Introduction
footnotetext: This work was supported by the National Social Science Foundation (No.24CYY107), the Fundamental Research Funds for the Central Universities (No.2024TD001), and the National Natural Science Foundation of China (No. 62576120).
@@
Although Large Language Models (LLMs) have demonstrated impressive multilingual capacities, there is still significant space for improving the performance on low-resource languages Huang et al. (2024); Al Nazi et al. (2024). To address this issue, especially avoiding costly post-training Muller et al. (2021); Yong et al. (2023), it is critical to fully investigate the multilingual understanding and transferring ability in LLMs.
@@
Recent research has increasingly focused on data-efficient methods, particularly Cross-lingual In-Context Learning (X-ICL) Winata et al. (2021); Tanwar et al. (2023); Al Nazi et al. (2024); Cahyawijaya et al. (2024), which surprisingly works well on low-resource languages, likely because LLMs are in-context low-resource language learners Brown et al. (2020b); Cahyawijaya et al. (2024). For instance, in the Arabic-to-Hebrew Bilingual Lexicon Induction (BLI) task, the zero-shot baseline accuracy in LLaMA 3 is 47.0%. However, simply specifying English as a bridge language in a zero-shot setting boosts accuracy to 64.5%, which even significantly outperforms the two-shot X-ICL. This observation motivates us to further explore: How can we improve cross-lingual capabilities of LLMs on low-resource languages by selecting an optimal bridge language in X-ICL? Should the selection be purely data-driven, favoring high-resource bridge languages Vulic et al. (2020)? Or can human linguistic knowledge, such as language genealogy, or established evolutionary taxonomies, offer a more effective alternative Stanczak et al. (2022); Wang et al. (2024)?
@@
To systematically investigate this issue, we leverage linguistic neurons Tang et al. (2024) that handle language features to guide optimal bridge language selection in X-ICL. However, there are two limitations when applying neuron-based interpretation Cao et al. (2024); Tang et al. (2024); Liu et al. (2024) on low-resource languages:
@@
$\bullet$ Inaccurate neuron activation. Current work often relies on multilingual corpora like Wikipedia Foundation (2024) to probe internal neurons, without verifying whether LLMs truly understand the multilingual input. This may lead to unreliable neuron activations, particularly for low-resource languages. When LLMs poorly understand the probe input, they may instead activate neurons for processing unfamiliar or noisy input.
@@
$\bullet$ Lacking guidance for cross-lingual transfer. Recent work argues that language-specific neurons do not facilitate cross-lingual transfer Mondal et al. (2025). This raises a critical question: if language-specific neurons cannot, can sharing neurons improve cross-lingual transfer in LLMs? This exploration is also important for transferring language neuron research to actionable strategies to enhance the multilinguality of LLMs.
@@
Motivated by this, we propose a simple yet effective bridge method, BridgeX-ICL, to improve LLMs’ cross-lingual capabilities, especially on low-resource languages. To address the inaccurate activation issue, we construct probe data by leveraging the ground-truth bilingual lexicon MUSE Conneau et al. (2017). We collect bilingual word pairs from MUSE that LLMs can translate accurately and use them to prompt the models for bidirectional translations, generating answers in both language directions. To address the cross-lingual guidance issue, we first explore overlap neurons’ features and their impact on cross-lingual transfer, and then propose a bridge selection strategy based on the Hilbert-Schmidt Independence Criterion (HSIC) Gretton et al. (2005). Furthermore, we measure the linguistic spectrum in LLMs based on overlapping-neurons and compare it with human language genealogy from Glottolog Trees Hammarström et al. (2023). We conduct extensive experiments on $4$ cross-lingual tasks and $15$ language pairs from $7$ diverse families. Our main contributions and findings are as follows:
@@
  • We construct accurate neuron probe data and use it to fully activate the anchored overlap neurons. We also propose an HSIC-based metric to quantify the similarity between overlapping neurons and specific neurons for optimal bridge selection in X-ICL.
  • We validate the generalization of BridgeX-ICL on $4$ cross-lingual tasks and $15$ language pairs. Here are empirical findings: 1) Strong neural overlaps align with human linguistic taxonomy within language families, but do not consistently hold across families. 2) Overlapping neurons embody shared semantic information, regardless of the language within or between families. 3) BridgeX-ICL improves the performance on cross-lingual tasks of BLI and MRC across 15 language pairs by an average of $6.02\%$ and $5.25\%$ over zero-shot baselines. 4) English is selected as the optimal bridge in 9 out of 15 language pairs in LLaMA (7 out of 15 in Mistral), indicating that high-resource, Latin-script languages tend to be the default bridge. We also find that non-Latin script languages like Chinese also show potential as effective bridges.
@@ 2.1 Cross-lingual In-context Learning
LLMs face significant challenges when applied to low-resource languages Costa-jussà et al. (2022); Muennighoff et al. (2023); Huang et al. (2024), mainly due to insufficient training data and the curse of multilinguality Conneau et al. (2020). To address these issues without updating model parameters, Cross-lingual In-context Learning (X-ICL), an extension of in-context learning (ICL), has recently gained attention Brown et al. (2020a). Prior studies Winata et al. (2021); Tanwar et al. (2023); Al Nazi et al. (2024); Cahyawijaya et al. (2024) have demonstrated that LLMs act as effective few-shot multilingual learners, with few-shot ICL even outperforming fine-tuned language-specific models on several tasks Winata et al. (2021). However, few-shot X-ICL’s performance is highly dependent on the context and the selection of examples, especially for unconventional or ambiguous languages Philippy et al. (2023); Al Nazi et al. (2024). Consequently, existing research mainly focuses on optimizing few-shot example selection. To the best of our knowledge, we are the first to explore X-ICL explicitly from the perspective of leveraging language bridges.
@@ 2.2 Linguistic Neuron in LLMs
Recent research Stanczak et al. (2022); Tang et al. (2024); Liu et al. (2024); Cao et al. (2024); Wang et al. (2024) has revealed that language-related neurons exist in FFN layers of transformer architecture. Deactivating these neurons will have a vital impact on LLMs’ multilingual capacities. Beyond uncovering multilingual mechanisms, some research has gone to explore the neuron pattern across languages Wang et al. (2024); Stanczak et al. (2022) and its impact on cross-lingual performance Mondal et al. (2025); Zhang et al. (2025). Specifically, Wang et al. (2024) observed that similar languages may not exhibit significant neuron sharing in LLMs like BLOOM, suggesting that neuron sharing does not fully align with language similarity. Furthermore, recent work argues that language-specific neurons do not facilitate cross-lingual transfer Mondal et al. (2025). This raises a critical question: whether sharing neurons can improve cross-lingual performance in LLMs. Motivated by these findings, we aim to further investigate LLM-internal neuron sharing across languages and its impact. In particular, we define a subset of language-overlapping neurons and explore whether they can serve as internal bridges to support cross-lingual inference.
@@ 3.1 Task Statement
Given a set of languages $\mathcal{L}=\{L_{1},...L_{|\mathcal{L}|}\}$, this work aims to measure the linguistic genealogy implicitly learned by LLMs from language overlapping neurons, then use the quantified linguistic similarity to guide the bridge language selection in X-ICL.
Figure 1 depicts three main steps of our approach: ① neuron probe data construction; ② language neurons and their overlapping detection; ③ bridge language selection, guided by the observed pattern of overlapping neurons and a modified HSIC dependency estimation, which selects the optimal $L_{*}$ from the candidate set $\mathcal{L}_{\text{candidate}}$ to facilitate X-ICL from a source language $L_{s}$ to a target language $L_{t}$.
@@ 3.2 Probe Data Construction
We employ two types of probe data for different purposes of language neurons identification and optimal bridge selection. The former is task independent, targeting language neurons, and can use existing multilingual corpora. In our work, we adopt FLORES+ NLLB Team et al. (2024), a high-quality parallel corpus released by Meta, and combine its development set and test set to obtain $2,000$ parallel sentences for each language.
Bridge selection for X-ICL needs to consider both language neurons and those that contribute to cross-lingual tasks. Inspired by findings on task-specific neurons Song et al. (2024), we propose that certain neurons directly influence cross-lingual transfer, and their manipulation and measurement should not rely solely on monolingual corpora. Therefore, we construct probe data by leveraging bilingual word translations as follows.
Probe Data Design. We collect $d$ (i.e., $100$) word pairs that LLMs can accurately translate. These word pairs are fed into the LLMs in both directions of $L_{1}\to L_{2}$ and $L_{2}\to L_{1}$, ensuring neurons linked to $L_{1}$ and $L_{2}$ are fully activated. Instead of feeding word pairs directly, we prompt LLMs to generate translations, which guarantees accurate neuron activation. Examples of probe data for $3$ language pairs are shown below.
@@ 3.3.1 Neurons in LLMs
Neuron identification follows Tang et al. (2024), which assumes that language neurons are mainly located in the Feed-Forward Network (FFN) layers. Given the transformation at the $i$-th layer:
@@
where $\tilde{\boldsymbol{h}_{i}}$ is the hidden state input to the $i$-th layer and $\sigma(\cdot)$ denotes the activation function. $\boldsymbol{W}_{1}^{i}\in\mathbb{R}^{d\times N}$ and $\boldsymbol{W}_{2}^{i}\in\mathbb{R}^{N\times d}$ are the learned parameters. Here, a neuron is defined as a linear transformation of a single column in $\boldsymbol{W}_{1}^{i}$ and there are $N$ neurons in each layer. The activation value of the $j$-th neuron is $\sigma(\tilde{\boldsymbol{h}_{i}}\boldsymbol{W}_{1}^{i})_{j}$. If this value exceeds 0, the neuron is considered an activated neuron.
@@ 3.3.2 Overlap Neuron Identification
First, we identify neurons $\mathcal{T}_{k}$ associated with each language $L_{k}$. Unlike existing work Wang et al. (2024); Mondal et al. (2025) using Language Activation Probability Entropy (LAPE) Tang et al. (2024) to identify neurons with high activation probability for one language but low for others, which is less effective to capture neuron relationships across languages, we identify neurons set $\mathcal{T}_{k}$ based on activation frequency. Let $f_{k,j}$ denote the activation frequency of neuron $n_{j}$ when processing tokens from language $L_{k}$. Neurons with the top $\tau\cdot N$ activation frequencies are selected into $\mathcal{T}_{k}$ based on a threshold $\tau$.
Overlap Neuron Definition. For languages $L_{u}$ and $L_{v}$, with associated neuron sets $\mathcal{T}_{u}$ and $\mathcal{T}_{v}$, the overlap neurons are defined as the interaction of $\mathcal{T}_{u}$ and $\mathcal{T}_{v}$. At the $i$-th FFN layer, we have $\mathcal{T}_{u,v}(i)=\mathcal{T}_{u}(i)\cap\mathcal{T}_{v}(i)$.
Linguistic Similarity Calculation. The linguistic similarity between $L_{u}$ and $L_{v}$ is quantified through the activation frequencies of their overlapping neurons. Let $\boldsymbol{f_{u}}=\{f_{u,1},f_{u,2},...,f_{u,|\mathcal{T}_{u,v}|}\}$ denote the activation frequency vector of neurons in $\mathcal{T}_{u,v}$ when processing tokens from $L_{u}$. The linguistic similarity between $L_{u}$ and $L_{v}$ is calculated:
@@
By computing pairwise similarities for all languages in $\mathcal{L}$, we obtain a comprehensive linguistic spectrum of the model.
3.3.3 Overlap Neuron Pattern
Second, we use the constructed probe data from FLORES+ to explore overlapping neurons’ features and their generalized impact on cross-lingual transfer so that we can utilize them to guide bridge language selection. We have two observations:
@@ 3.3.3 Overlap Neuron Pattern
Figure 3: Layer-wise latent embeddings projected with MDS in French-Swahili translation. A rainbow-colored path traces the latent embeddings across 32 layers. The predicted Swahili tokens are in green, and their correct English tokens are in orange.
$\bullet$ Similar languages share more neurons than distant ones. For example, Arabic-Hebrew within the same language family has more overlapping neurons than Arabic-Swahili across families, as presented in Figure 2. This suggests the potential of neural overlap to measure language distance.
$\bullet$ Overlap neurons are predominantly concentrated in the middle and final layers, serving distinct roles of semantic understanding and language decoding for next-token prediction. This neural function is further evidenced by neurons deactivation shown in Figure 7. Specifically, neurons in final layers are task-related and responsible for cross-lingual generation. To examine whether middle-layer neurons handle semantic understanding, we further employ a technique called logit lens Nostalgebraist (2020) to visualize the latent semantic embeddings across layers. We visualize French-Swahili translation using $60$ word pairs that LLaMA 3 translates accurately. We extract the model’s latent embeddings at each layer for next-token prediction and project them into a 2D space using classical multidimensional scaling (MDS), presented in Figure 3. The embedding trajectory is marked in a rainbow-colored path (e.g., red = layers 1-4, violet = layers 25-32). We can observe French inputs and their corresponding correct English next tokens cluster in middle layers, indicating that LLMs rely on the knowledge in high-resource languages like English to perform cross-lingual reasoning there. Neurons in middle layers should be prioritized over those in final layers when quantifying language similarity.